# Generalized Out-of-Distribution Detection and Beyond in Vision Language Model Era: A Survey

### v1

@@ 8 Future Directions
In this section, we discuss the future directions of OOD detection and UPD. For OOD detection, we explore not only OOD detection for VLMs but also single-modal OOD detection, with a specific focus on emerging challenges as VLMs evolve. For a discussion of the long-standing challenges in OOD detection, we can refer the readers to the previous generalized OOD detection paper [22].
8.1 OOD Detectin for Vision Language Models
@@
a. Hard OOD Detection  Hard OOD detection will become increasingly important in the future due to its high practicality and the challenging nature of the problem. Hard OOD detection utilizes the OSR benchmark setup, where some classes within a single dataset are designated as ID and others as OOD. In this field, not only small datasets such as ImageNet-10 and ImageNet-20 [26] but also datasets with a larger number of classes such as ImageNet-protocol [54] have been proposed. Many existing studies, such as LoCoOp [28] and LSN [44], primarily use the common ImageNet OOD benchmark, so hard OOD detection has not yet been well studied. This field will develop further in the future.
@@
b. Post-hoc Methods  To propose post-hoc methods is important for the fundamental performance improvement of CLIP-based OOD detection. The methods of directly employing an ID classifier such as MCM [26] are called post-hoc methods. Prior to CLIP, various approaches were proposed  [101, 102, 103, 104, 105, 183, 184, 108, 109, 111]. However, CLIP-based post-hoc methods often underperform methods with additional OOD prompts, so they are not extensively researched in zero-shot OOD detection. However, we should focus on the scalability of post-hoc methods. The post-hoc methods [26, 96] can be easily applied to many subsequent methods [28, 185, 44, 47], bringing fundamental performance improvements. Furthermore, very recently, post-hoc methods specifically tailored for prompt learning methods have also emerged [86]. Therefore, proposing post-hoc methods and demonstrating the improvements not only in zero-shot but also in subsequent few-shot settings [28, 185] is crucial, even if they underperform methods using OOD prompts in the zero-shot setting. This field should continue evolving, mirroring its growth before the advent of CLIP.
@@
c. Bridging the Gap with Closed-set Classifiers.  OOD detection ensures the safety of ID classifiers, so it is crucial to bridge the gap between the advancements in existing closed-set classifiers and OOD detection. Currently, the representative method for few-shot OOD detection is LoCoOp [28], a text prompt learning method based on CoOp. However, in closed-set settings, few-shot learning methods based on text prompt learning have been proposed other than CoOp [186, 187, 188]. Furthermore, text-based prompt learning methods only train the text prompts, so they cannot handle differences in image domains. Therefore, adopting methods that can handle image domain differences [189, 190, 191] for OOD detection is essential for bridging the gap with closed-set ID classifiers.
@@
d. Training-free Few-shot OOD Detection  The research direction of training-free few-shot OOD detection is still in its infant stage, with only one existing study [99]. Given the prevalence of training-based methods in few-shot OOD detection, proposing methods that do not require training is crucial. Considering the advancements of training-free methods in CLIP-based anomaly detection, we anticipate a similar trajectory for CLIP-based OOD detection. Future directions include refining adapter-based methods or leveraging external knowledge such as retrieval augmentation [192, 193]. Addressing training-free few-shot OOD detection is a pivotal step towards realizing more computationally efficient OOD detection in the future.
@@
e. Full-spectrum OOD Detection  CLIP-based full-spectrum OOD (FS-OOD) detection is a promising research area [115, 116]. In practical applications, there is a strong motivation to create models that can not only detect semantically shifted OOD inputs but also generalize to covariate-shifted data [116, 194]. Within CLIP-based methods, OOD detection and generalization are often discussed in separate contexts [28, 195], resulting in a trade-off between detection and generalization performance [98]. Furthermore, there exists potential ambiguity in the distinction between FS-OOD detection and open-set domain generalization. To eliminate the potential ambiguity, we have formulated a potential problem setting called hard FS-OOD detection. We hope this survey inspires further advancements and developments for FS-OOD detection.
f. Open-vocabulary OOD Detection  Open-vocabulary OOD (OV-OOD) detection has a high practical potential, but it is still in its infant stages [47]. In particular, as explained in Sec. 7.1, zero-shot OV-OOD detection is a potential research field. We hope that this survey paper will inspire future efforts in OV-OOD detection.
@@
g. Object-level OOD Detection  Object-level OOD detection remains an unexplored area. As discussed in Sec. 5.4, this is due to the vastness of the OOD space, which makes it difficult to identify OOD objects using texts. To pave the way for future advancements in object-level OOD detection/segmentation, foundation models for localization, such as SAM [57], offer a promising solution. By integrating these models with methods such as MCM [26], we can potentially achieve object-level OOD detection and segmentation, opening up a new frontier in OOD detection research.
@@ 8.2 Single-modal OOD Detection
a. Leveraging Large Pre-trained Models  Leveraging large pre-trained models is important for single-modal OOD detection. Numerous methods for OOD detection conduct experiments using backbones trained from scratch and do not utilize pre-trained models [22, 119, 101, 196, 104, 197, 108, 109]. In a recent study, Miyai et al. [198] systematically investigated the impact of pre-training on OOD detection from both the perspectives of the types of OOD data and pre-training algorithms [199, 200]. Dong et al. [201] explored parameter-efficient learning for single-modal OOD detection and proposed DSGF, which leverages both fine-tuned features and original pre-trained features. While leveraging large pre-trained models [202] with lightweight tuning is an active area of research in single-modal closed-set classification [191, 203], there have been limited studies for single-modal OOD detection, which presents a promising avenue for future research.
b. Real-world Benchmarks and Evaluations  Considering the future development of CLIP-based OOD detection, there should be increasing focus on expanding the scope of benchmarks to encompass real-world scenarios where CLIP is less applicable. For instance, recently, Baek et al. [204] introduced ImageNet-ES, consisting of variations in environmental and camera sensor factors by directly capturing 202k images with an actual camera in a controlled testbed, which bridges the gap between the common benchmark and the real-world scenario. Besides, utilizing datasets such as WILDS [205, 206], which consider real-world data shifts, or datasets for medical OOD detection [207], can provide valuable insights, especially in safety-critical applications like autonomous driving and medical image analysis.
8.3 Unsolvable Problem Detection
a. Exploring Effective Solutions  It is important to propose effective solutions for UPD. One of the potential approaches is to adapt the methodologies of OOD detection to UPD. For example, the perplexity of the LVLM’s response could be used as a score to identify unsolvable queries. Furthermore, proposing model-agnostic post-hoc methods is also important to enhance the reliability of many LVLMs. Therefore, incorporating the concepts of OOD detection techniques for UPD is an important direction for future work.
b. Extension to Diverse Benchmarks  MM-UPD Bench consists of general QA datasets. However, UPD can be incorporated into more diverse benchmarks including domain-specific knowledge for advanced reasoning [208, 163] and multi-image understanding [164]. For example, MuirBench [164] incorporated the concept of UPD as a metric for robust evaluation. Integrating the concept of UPD into benchmarks is essential for evaluating the robustness and trustworthiness of LVLMs in their target tasks.
c. Theoretical Understanding of UPD  Theoretically understanding the reasons behind the difficulty of UPD could provide the community with valuable insights. Theorizing the behavior of LVLMs poses a shared challenge in the field, highlighting the importance of collaborative efforts within the community.
@@ 9 Conclusion
In this survey, we comprehensively review the evolution of the five problems including AD, ND, OSR, OOD detection, and OD in the VLM era, and propose a framework of generalized OOD detection v2. Our framework identifies OOD detection and AD as the primary challenges in the VLM era, which highlights the demanding challenges and fosters collaborative efforts among each community. By articulating the shifts in the definitions, problem settings, and benchmarks, we encourage subsequent works to accurately understand their evolving target problems in the VLM era. By sorting out the methodologies, we hope that readers can easily grasp the mainstream methods, identify important baselines and novel problem settings, and propose future solutions. By shedding light on recent studies in the LVLM era, we hope that researchers within each community can identify promising research directions in this emerging era. By providing future directions, we hope that our survey will clarify the tasks to be tackled by future works in the VLM and LVLM era, thereby facilitating future advances in the right direction.
@@ Acknowledgment
This work was supported by JST BOOST, Japan Grant Number JPMJBS2418 and JST JPMJCR22U4. We thank Toyooka Mashiro (AYM Lab at UTokyo) for his valuable assistance in designing the figures, and Kazuki Egashira, Yuki Imajuku, Takubon Son, and Zaiying Zhao (AYM Lab at UTokyo) for valuable feedback on the paper.

### v2

@@ 8 Future Directions
In this section, we discuss the future directions of OOD detection. We explore not only OOD detection for VLMs but also single-modal OOD detection, with a specific focus on emerging challenges as VLMs evolve. For a discussion of the long-standing challenges in OOD detection, we can refer the readers to the previous generalized OOD detection paper (Yang et al., 2024).
8.1 OOD Detection for Vision Language Models
@@
a. Hard OOD Detection  Hard OOD detection will become increasingly important in the future due to its high practicality and the challenging nature of the problem. Hard OOD detection utilizes the OSR benchmark setup, where some classes within a single dataset are designated as ID and others as OOD. In this field, not only small datasets such as ImageNet-10 and ImageNet-20 (Ming et al., 2022a) but also datasets with a larger number of classes such as ImageNet-protocol (Palechor et al., 2023) and ImageNet-X (Noda et al., 2025) have been recently proposed. Many existing studies, such as LoCoOp (Miyai et al., 2023b) and LSN (Nie et al., 2023), primarily use the common ImageNet OOD benchmark, so hard OOD detection has not yet been well studied. This field will develop further in the future.
@@
b. Post-hoc Methods  To propose post-hoc methods is important for the fundamental performance improvement of VLM-based OOD detection. The methods of directly employing an ID classifier such as MCM (Ming et al., 2022a) are called post-hoc methods. Prior to CLIP, various approaches were proposed  (Hendrycks & Gimpel, 2017; Liang et al., 2018; Lee et al., 2018; Liu et al., 2020; Sastry & Oore, 2020; Sun & Li, 2022; Sun et al., 2022; Sastry & Oore, 2019). Nevertheless, VLM-based post-hoc methods often underperform methods with additional OOD prompts, so they are not extensively researched in zero-shot OOD detection. However, we should focus on the scalability of post-hoc methods. The post-hoc methods (Ming et al., 2022a; Miyai et al., 2025b) can be easily applied to many subsequent methods (Miyai et al., 2023b; Ming et al., 2022b; Nie et al., 2023; Li et al., 2024b), bringing fundamental performance improvements even for the few-shot setting. Furthermore, recently, post-hoc methods specifically tailored for prompt learning methods have also emerged (Jung et al., 2024). Therefore, proposing post-hoc methods and demonstrating the improvements not only in zero-shot but also in subsequent few-shot settings (Miyai et al., 2023b; Ming et al., 2022b) is crucial. This field should continue evolving, mirroring its growth before the advent of CLIP.
@@
c. Bridging the Gap with Closed-set Classifiers.  OOD detection ensures the safety of ID classifiers, so it is crucial to bridge the gap between the advancements in existing closed-set classifiers and OOD detection. Currently, the representative method for few-shot OOD detection is LoCoOp (Miyai et al., 2023b), a text prompt learning method based on CoOp. However, in closed-set settings, VLM-based few-shot learning methods have been proposed other than CoOp (Chen et al., 2023a; Bulat & Tzimiropoulos, 2023; Lu et al., 2022). Therefore, adopting recent methods for OOD detection is essential for bridging the gap with closed-set ID classifiers.
@@
d. Training-free Few-shot OOD Detection  The research direction of training-free few-shot OOD detection is still in its infant stage (Chen et al., 2024b; Zhang et al., 2024f). Considering the advancements of training-free methods in VLM-based AD, we anticipate a similar trajectory for VLM-based OOD detection. Future directions include refining adapter-based methods or leveraging external knowledge such as retrieval augmentation (Udandarao et al., 2023; Ming & Li, 2024b). Addressing training-free few-shot OOD detection is a pivotal step towards realizing more computationally efficient OOD detection in the future.
@@
e. Full-spectrum OOD Detection  VLM-based full-spectrum OOD (FS-OOD) detection is a promising research area (Lu et al., 2023; Yang et al., 2023). In practical applications, there is a strong motivation to create models that can not only detect semantically shifted OOD inputs but also generalize to covariate-shifted data (Yang et al., 2023; Bai et al., 2023). Within VLM-based methods, OOD detection and generalization are often discussed in separate contexts (Miyai et al., 2023b; Khattak et al., 2023), resulting in a trade-off between detection and generalization performance (Lafon et al., 2024). Therefore, we need to discuss both aspects together to mitigate the trade-off.
f. Open-vocabulary OOD Detection  Open-vocabulary OOD (OV-OOD) detection has a high practical potential, but it is still in its infant stages (Li et al., 2024b). The open-vocabulary setting has been actively explored in AD ( 5.1) and has the potential to become increasingly important in the field of OOD detection as well.
@@
g. Object-level OOD Detection  Object-level OOD detection remains an unexplored area in VLM-based OOD detection. As discussed in Sec. 5.4, this is due to the vastness of the OOD space, which makes it difficult to identify OOD objects using texts. To pave the way for future advancements in object-level OOD detection/segmentation, foundation models for localization, such as SAM (Kirillov et al., 2023), offer a promising solution. By integrating these models with methods such as MCM (Ming et al., 2022a), we can potentially achieve object-level OOD detection and segmentation, opening up a new frontier in OOD detection research.
@@ 8.2 Single-modal OOD Detection
a. Leveraging Large Pre-trained Models  Leveraging large pre-trained models is important for single-modal OOD detection. Numerous methods for OOD detection conduct experiments using backbones trained from scratch and do not utilize pre-trained models (Yang et al., 2024; Zhang et al., 2024b; Hendrycks & Gimpel, 2017; Leys et al., 2018; Liu et al., 2020; Wang et al., 2022a; Sun & Li, 2022; Sun et al., 2022). In a recent study, Miyai et al. (2023a) systematically investigated the impact of pre-training on OOD detection from both the perspectives of the types of OOD data and pre-training algorithms (Chen et al., 2020; Caron et al., 2021). Dong et al. (2023) explored parameter-efficient learning for single-modal OOD detection and proposed DSGF, which leverages both fine-tuned features and original pre-trained features. While leveraging large pre-trained models (Dosovitskiy et al., 2021) with lightweight tuning is an active area of research in single-modal closed-set classification (Hu et al., 2022; Zhang et al., 2024e), there have been limited studies for single-modal OOD detection, which presents a promising avenue for future research.
b. Real-world Benchmarks and Evaluations  Considering the future development of VLM-based OOD detection, there should be increasing focus on expanding the scope of benchmarks to encompass real-world scenarios where CLIP is less applicable. For instance, recently, Baek et al. (2024) introduced ImageNet-ES, consisting of variations in environmental and camera sensor factors. Besides, utilizing datasets such as WILDS (Koh et al., 2021; Cultrera et al., 2023), which consider real-world data shifts, or datasets for medical OOD detection (Hong et al., 2024), can provide valuable insights, especially in safety-critical applications such as autonomous driving and medical image analysis.
@@ 9 Conclusion
In this survey, we comprehensively review the evolution of the five problems including AD, ND, OSR, OOD detection, and OD in the VLM era, and propose a framework of generalized OOD detection v2. Our framework identifies OOD detection and AD as the primary challenges in the VLM era. By articulating the shifts in the definitions, problem settings, benchmarks and methodologies, we encourage subsequent works to accurately understand their evolving target problems in the VLM era. By shedding light on recent studies in the LVLM era, we hope that researchers within each community can identify promising research directions in this emerging era. By providing future directions, we hope that our survey will clarify the tasks to be tackled by future works in the VLM era, facilitating future advances in the right direction.
@@ Acknowledgment
This work was supported by JST BOOST, Japan Grant Number JPMJBS2418 and JST JPMJCR22U4. We thank Toyooka Mashiro (AYM Lab at UTokyo) for his valuable assistance in designing the figures, Kazuki Egashira, Yuki Imajuku, Takubon Son, and Zaiying Zhao (AYM Lab at UTokyo) for valuable feedback on the paper, and Shiho Noda for helping experiments.